OpenStudy (anonymous):

Linear functions unit portfolio. Write an equation in slope-intercept, point-slope, or standard form. Explain why you chose the form you did. This is about cab services and fees. The cabs flat rate is $3.00 and plus $0.30 per mile. Help Please

OpenStudy (mrmanman305):

well they give you a -intercept and a slope

OpenStudy (mrmanman305):

which one of those forms need that information.

OpenStudy (anonymous):

I think it would be Standard form.

OpenStudy (mrmanman305):

why do you think that

OpenStudy (fortytherapper):

It's saying you can write it in any form, but you have to use given information to write one

OpenStudy (mrmanman305):

which one would be the simplest one though

OpenStudy (fortytherapper):

Standard form by far according to this scenario

OpenStudy (mrmanman305):

I would say slope intercept

OpenStudy (fortytherapper):

Standard form being A(x) + B(x) = C

OpenStudy (fortytherapper):

You would just have to plug in the given values

OpenStudy (anonymous):

y = 3.00 + 0.30x would that equation work for any of them. Idk what any of those forms are tbh.

OpenStudy (fortytherapper):

Point slope is a good one too

OpenStudy (mrmanman305):

y=mx+b y=0.30x+3.00

OpenStudy (fortytherapper):

Yeah @Jay_me2901 , you used slope intercept. Good job

OpenStudy (fortytherapper):

As Mrman said y = mx+b

OpenStudy (anonymous):

Thank you both. If you stick around ill probably need more help on some other questions. Its a big portfolio project

OpenStudy (mrmanman305):

Will do